Physical Chemistry

The values of KSp of CaCO3 and CaC2O4 are 4.7x10-9 and 1.3 x10-9 respectively at 25degree centigrade .If the mixture of these two is washed with water ,what is the concentration of Ca+2 ions in water

praveen kumar
11 Years agoGrade

1 Answer

Naveen Kumar
11 Years ago
Dear student also specify the amount of water added.
Suppose we have solubility of these compounds x and y respectively.
CaCO3................>Ca2+ + CO32-
…............................x.............x
x*x=4.7*10^-9, x=6.7*10^-5

CaC2O4..................> Ca2+ + C2O42-
….................................y.............y
y*y=1.3*10^-9, y=3.6*10^-5
Now if both are mixed, then Ca2+ would be the common.